# Sharp transitions in small exciton spectra for multi-orbital lattice systems

### v2

@@ IV.1 The range of validity of the continuum approximation
One way to validate the method described above is to verify that in the limit $U\rightarrow 0$, its results agree with those predicted by the continuum approximation. This will demonstrate that our method is also suitable to calculate the spectra of large excitons, although in this case the above-mentioned cutoff is large (it has to exceed the exciton radius) and the computation becomes much more expensive than for small excitons.
Figure 1: 1D exciton with K=0{K}=0 binding energy (left panel) and its radius in units $a=1$ (right panel) from the continuum approximation (blue line) and a 1D lattice model (red circles) with a one-orbital conduction band with $t_{c}=1$, and a $s,p_{x}$ two-orbital valence band with $t_{ss}=1.1,t_{pp}=-1,t_{sp}=0.5$. The corresponding valence band structure is shown in Fig. 2. Here, we assume $U_{s}=U_{p}=U$. The vertical dashed line marks the value of $U$ above which the continuum approximation starts to fail. See text for more details.
Figure 2: The two valence bands $E_{v}({k,\pm})$ (red lines) corresponding to the parameters used in Fig. 1. The top of the valence band is at $k=0$, as is the bottom of the conduction band (not shown), indicating a direct gap $K=0$. The dashed blue line shows the continuum approximation, which matches the exact dispersion to within $1\%$ for $|k|\lesssim\pi/6$ (dashed vertical lines).
Figure 1 compares the exciton binding energy for momentum K=0{K}=0, and the corresponding exciton radius $\xi$ as predicted by a 1D continuum model (blue lines; the relevant equations are listed in Appendix D) and a 1D lattice model with a $s,p_{x}$ two-orbital valence band (solid red circles). The magnitudes of the hopping integrals are $t_{c}=1$, $t_{ss}=1.1,t_{pp}=-1$ and $t_{sp}=0.5$ (the signs due to the appropriate orbital overlaps are included). For these values, the semiconductor has a direct gap, which is why we consider ${\bf K}=0$. This is demonstrated in Fig. 2, which shows the dispersions of the two valence bands $E_{v}({k},\pm)$ in the Brillouin zone (red lines).
For this specific model, the continuum approximation is accurate for $U\leq U_{c}\approx 2.2t_{c}$, after which it becomes increasingly worse, see Fig. 1. This corresponds to an exciton radius $\xi\gtrsim 2a$, which at first sight is a surprisingly low bound for a 'large' exciton. However, we can understand this as follows. For a given $\xi$, top-valence and conduction band states with momenta up to $k\sim 1/\xi$ contribute the most to the $K=0$ exciton $|K=0,{\rm exc}\rangle=\sum_{k}\phi_{k}c^{\dagger}_{k}v_{k,+}|{\rm GS}\rangle$. Here, $\xi\gtrsim 2a$ implies $|k|\lesssim{1\over 2a}\approx{\pi\over 6a}$. This value is shown by the vertical dashed lines in Fig. 2, and indeed provides a good bound for the top-valence band states described well by the continuum approximation (blue parabola).
@@
This simple procedure provides a practical recipe for estimating whether a continuum approximation is reasonable when the dispersions of the conduction and valence are known. As we point out in the next section, it may still be very misleading in a multi-orbital system.
@@ IV.2 Qualitative failure of the continuum approximation in multi-orbital systems
Different Coulomb attractions for holes located in different orbitals add a second important ingredient in deciding the exciton spectra, besides the dispersion of the top valence band in the multi-orbital system, discussed above. This is because a stronger Coulomb attraction for one of the orbitals may lead to a larger binding energy for excitons involving valence band states which has a maximal character of that type, even if these states are not at the top of the valence band.
@@
Such an example is exemplified in Fig. 3. The solid black line shows $E_{\rm gap}(K)$, i.e. the lower edge of the electron-hole continuum, for the same parameters as in Fig. 1. Clearly, this is a direct-gap semiconductor with the minimum gap at $K=0$. Accordingly, the continuum approximation always predicts that the maximum exciton binding energy must appear at $K=0$. The exciton spectrum of the 1D lattice agrees with this prediction if $U_{s}=U_{p}=0.5$, shown by the red circles. Indeed, here the strength of the attraction is the same independent of the character of the top-valence band, and therefore the most stable exciton appears where the electron-hole excitation gap is minimum. However, if we choose $U_{p}=3U_{s}=1.5$ (blue crosses), then the exciton minimum moves to $K=\pm\pi$ because the top-valence band has a predominant $p$-character near $k=\pi$ (for this model). Note that the exciton dispersion near $K=0$ is little affected by the increase of $U_{p}$, consistent with the fact that near $k=0$ the top-valence band has primarily $s$-character in our model.
To summarize, if the continuum approximation simplifies the multi-orbital valence bands' dispersion to a single parabola centered at the top of the top-valence band, it can never reproduce the phenomenology exemplified in this subsection, because information regarding the orbital character of the actual valence band is lost. Of course, one could proceed to a 'multi-valley' type of continuum approximation that would do better, however, that removes the top advantage of the continuum approximation, namely its ability to produce analytical results. If numerical computations are needed, studying a realistic lattice model becomes more efficient, beside being accurate when describing small excitons.
@@
As an example, we use the 2D triangular lattice with the three $d$ orbital bases introduced in Section II. Symmetry guarantees that their respective overlaps with an $s$-type conduction band orbital are the same, hence we set $U_{\alpha}=U$ for all three valence orbitals. (A short discussion of the effect of unequal $U_{\alpha}$ can be found in Appendix C). The other parameters are listed in Appendix A.
@@
Figure 4 shows the spectral weight $A_{\alpha\alpha}({\bf K},\omega)=-{1\over\pi}\text{Im}G_{\alpha\alpha}({\bf K},\bm{0},z)$ for $\bf K$} along high-symmetry lines in the Brillouin zone, for $\alpha=d_{x^{2}-y^{2}}$ (top panel), $\alpha=d_{xy}$ (middle panel) and $\alpha=d_{3z^{2}-r^{2}}$ (bottom panel). The thick black line marks the lower edge of the electron-hole continuum, $E_{\rm gap}({\bf K})$. Above this edge, the spectral function contains continuum states, but on the color scale used, it is too weak to be clearly visible. The interesting features are the three exciton bands visible well inside the gap (due to the considerable $U$ value used). In the $d_{3z^{2}-r^{2}}$ projection (bottom panel) most of the intensity resides in the highest-energy branch and is therefore less relevant to the ground-state discussion. Along the $\Gamma-M-K$ cuts, the lowest branch appears with comparable intensity in the top two panels which show that it has comparable amounts of $d_{x^{2}-y^{2}}$ and $d_{xy}$ character for these momenta. By contrast, along the $K-\Gamma$ direction, top panel shows two exciton branches while the middle panel shows a third, different branch. Clearly, along this cut there is no mixing between these two orbitals, and the excitons have either $d_{x^{2}-y^{2}}$ or $d_{xy}$ character.
Figure 5: (top) Dispersion of the lowest-energy exciton $E_{\rm exc}({\bf K})$ for different values of $U$. The blue dashed line is the exciton energy calculated with a cut-off at $\delta=2$ (see Appendix G for details), and the dot marks its predicted lowest-energy exciton; (bottom) $E_{\rm exc}({\bf K})$ when ${\bf K}=M$ (blue triangles) and when $\bf K$} is the indirect gap, i.e. local minimum for $E_{\rm gap}({\bf K})$ along the $K-\Gamma$ line (red circles).
@@
The continuum approximation would predict that the lowest energy exciton appears along the $K-\Gamma$ line, at the momentum $\bf K$} for which $E_{\rm gap}({\bf K})$ has a minimum; this is the momentum of the indirect gap of this semiconductor model. The top panel indicates that the lowest energy exciton branch has indeed a minimum at that momentum, however a more careful inspection shows that the global minimum is actually at the $M$ point. We note that the $M$ point is not even a local minimum for $E_{\rm gap}({\bf K})$!
@@
To understand what happens, we plot in the top panel of Fig. 5 the energy of the lowest-energy exciton branch along the same cuts, for increasing values of $U$. As expected, the exciton moves to lower energies as the attraction $U$ increases. Its dispersion follows closely the gap energy $E_{\rm gap}({\bf K})$ for small values of $U$, however, this changes as $U$ increases.
Figure 7: Probability $|\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle|^{2}$ plotted as a function of the distance $\bm{\delta}$ between electron and hole, for the lowest-energy exciton and for $\bf K$} equal to the indirect gap momentum, for $U=12$. The probabilities are normalized $\sum_{\beta,\bm{\delta}}|\langle{\bf K},{\rm exc}|\beta,{\bf K},\bm{\delta}\rangle|^{2}=1$. The projections are on orbital $\beta=d_{x^{2}-y^{2}}$ (top), $\beta=d_{xy}$ (middle) and $\beta=d_{3z^{2}-r^{2}}$ (bottom). The triangular lattice sites are shown by the white circles.
The bottom panel tracks the exciton energy at the momentum $M$ and at the indirect gap momentum (at minimum for $E_{\rm gap}({\bf K})$ along the $K-\Gamma$ line). The two curves cross at $U\approx 10$, marking a sharp transition in the momentum and character of the lowest-energy exciton. This is another example of a sharp transition that is impossible within the continuum approximation.
This change in the shape of the dispersion of the lowest-energy exciton band is illustrated graphically in Fig. 6, where the energy contour of $E_{\rm exc}({\bf K})$ is plotted in the Brillouin zone. For the smaller $U=2.5$ value (top panel) its minimum is clearly along the $\Gamma-K$ cut. However, as $U$ increases, the energy at the $M$ points drops faster such that for $U=12$ (bottom panel) the minimum is at the $M$ point.
@@
This transition can be attributed to the symmetry already mentioned, which prevents mixing of $d_{x^{2}-y^{2}}$ and $d_{xy}$ orbitals along the $\Gamma-K$ line. This is illustrated by the real-space probability plots to find the hole at a distance $\bm{\delta}$ from the electron in each of the three possible orbitals, shown in Fig. 7 for the indirect band-gap momentum and in Fig. 8 at the $M$ point.
First, both of these excitons are Frenkel-like, with most of the probability concentrated at $\bm{\delta}=0$. The exceptions are the $d_{x^{2}-y^{2}}$ and the $d_{3z^{2}-r^{2}}$ orbitals (top and bottom panel in Fig. 7, respectively) where symmetry prevents weight at $\bm{\delta}=0$ and instead there is a very small amount of weight at $|{\bm{\delta}}|=a$. This is consistent with the vanishing spectral weight for this exciton in the top plot of Fig. 4. By contrast, the $M$ point exciton has considerable on-site probability in all orbitals, see Fig. 8).
[…]
@@ V Conclusions
In this work, we adapt a well-established approach for studying interacting few-fermion systems to the study of exciton spectra in multi-orbital lattice models. This method is particularly efficient for small excitons, such as those known to appear in a variety of semiconductors.
The results of the lattice models agree with the predictions of the continuum approximation in the limit of weak electron-hole attraction. This comparison allowed us to propose a simple criterion to estimate when the continuum approximation is likely to be accurate.
However, we also demonstrated that multi-orbital models exhibit exciton physics that is impossible to anticipate with a simple continuum approximation. Specifically, we presented two simple examples in which the momentum of the lowest-energy exciton does not match the momentum of the gap in certain regions of the parameter space. These examples only scratch the surface in terms of possible new behavior of small excitons in complex, multi-orbital bands, but help to illustrate the need to use realistic lattice models when modeling small excitons in specific semiconductors.

### v3

@@ IV.1 The range of validity of the continuum approximation
One way to validate the method described above is to verify that in the limit $U\rightarrow 0$, its results agree with those predicted by the continuum approximation. This will demonstrate that our method is also suitable for calculating the spectra of large excitons, although in this case the above-mentioned cutoff is large (it has to exceed the exciton radius) and the computation becomes expensive.
Figure 2: The conduction band (orange dotted line) and the two valence bands $E_{v}({k,\pm})$ (symbols) for the parameters listed in Fig. 1. The top/bottom of the valence/conduction band is at $k=0$, indicating a $K=0$ direct gap. The dashed black line shows the continuum approximation for the valence band, which matches the exact dispersion to within $1\%$ for $|k|\lesssim\pi/6$ (dashed vertical lines). The blue circles indicate the weight of $p_{x}$ orbital character while the red squares indicate the weight of $s$ orbital character.
Figure 1 compares the binding energy of the exciton with momentum K=0{K}=0 and its corresponding exciton radius $\xi$ as predicted by the 1D lattice model with a $s,p_{x}$ two-orbital valence band (solid red circles) and by its corresponding continuum approximation (blue lines; the relevant equations are listed in Appendix D). For the lattice model, the hopping integrals are set to $t_{c}=1$, $t_{ss}=1.1,t_{pp}=-1$ and $|t_{sp}|=0.5$ (the signs due to the orbital overlaps are included in the Hamiltonian). For these values, the semiconductor has a direct gap, which is why we focus on the lowest-energy exciton with K=0{K}=0. The direct gap is demonstrated in Fig. 2, which shows the dispersion of the conduction band $E_{c}(k)=-2t_{c}\cos(k)$ (dotted orange line) and of the two valence bands $E_{v}({k},\pm)$ in the Brillouin zone (symbols illustrate their orbital character, with blue/red associated with the $p_{x}$ and $s$ orbitals, respectively). The top of the top-valence band is at $k=0$.
For these parameters, Fig. 1 shows that the continuum approximation is accurate for $U\leq U_{c}\approx 2.2t_{c}$, after which it becomes increasingly worse, see Fig. 1. This corresponds to an exciton radius $\xi\gtrsim 2a$, which at first sight is a surprisingly low bound for a 'large' exciton. However, we can understand this as follows. For a given $\xi$, top-valence and conduction band states with momenta up to $k\sim 1/\xi$ contribute the most to the $K=0$ exciton $|K=0,{\rm exc}\rangle=\sum_{k}\phi_{k}c^{\dagger}_{k}v_{k,+}|{\rm GS}\rangle$. Here, $\xi\gtrsim 2a$ implies $|k|\lesssim{1\over 2a}\approx{\pi\over 6a}$. This value is shown by the vertical dashed lines in Fig. 2, and indeed provides a good bound for the region where the top-valence band states are described well by the continuum approximation (blue parabola).
@@
This simple procedure provides a practical recipe for estimating whether a continuum approximation is reasonable if the dispersions of the conduction and valence bands are known. As we point out in the next section, it may still be misleading in a multi-orbital system.
@@ IV.2 Qualitative failure of the continuum approximation in multi-orbital systems
Different Coulomb attractions for holes located in different orbitals add a second important ingredient in determining the exciton spectra, besides the dispersion of the top valence band in the multi-orbital system, discussed above. This is because a stronger Coulomb attraction for one of the orbitals may lead to a larger binding energy for excitons involving valence band states which have a maximal character of that type, even if these states are not at the top of the valence band.
@@
Such an example is illustrated in Fig. 3. The solid black line shows $E_{\rm gap}(K)$, i.e. the lower edge of the electron-hole continuum for different total momenta K{K}, for the same parameters as in Fig. 1. Clearly, this is a direct-gap semiconductor with the minimum gap at $K=0$. Accordingly, the simplest continuum approximation predicts that the maximum exciton binding energy must appear at $K=0$. The exciton spectrum of the 1D lattice agrees with this prediction if $U_{s}=U_{p}=0.5$, shown by the red circles. Indeed, here the strength of the attraction is the same independent of the character of the top-valence band, and therefore the lowest energy exciton appears where the electron-hole excitation gap is minimum. However, if we choose $U_{p}=3U_{s}=1.5$ (blue crosses), then the exciton minimum transitions to $K=\pm\pi$ because the top-valence band has a predominant $p$-character near $k=\pi$, see Fig. 2. Note that the exciton dispersion near $K=0$ is little affected by the increase of $U_{p}$, consistent with the fact that near $k=0$ the top-valence band has primarily $s$-character in our model.
To summarize, if the continuum approximation simplifies the multi-orbital valence bands' dispersion to a single parabola centered at the top of the top-valence band, it can never reproduce the sharp transition exemplified in this subsection, because information regarding the orbital character of the actual valence band is lost. Of course, for Wannier-like excitons, one could proceed to a ``multi-valley" type of continuum approximation that would do better, although that removes the top advantage of the continuum approximation, namely its ability to produce analytical results. For small, Frenkel-like excitons, however, the bands need to be accurately described in the full Brillouin zone, and this simply cannot be achieved quantitatively with any quadratic-type of approximation, whether single- or multi-valley. A multi-valley continuum approximation might still be made to be qualitatively accurate even in this case, although we think that is less likely for the 2D example discussed next. Nevertheless, we believe that our method makes it easier to compute the exact spectra of small, Frenkel-like excitons than it would be to develop and validate a continuum approximation whose predictions might remain qualitatively correct in this limit.
@@
As an example, we use the 2D triangular lattice with the three $d$ orbital bases introduced in Section II. Figure 4 shows the conduction band (blue solid line) and the three valence bands and their orbital character. Specifically, blue circles indicate the weight of the $d_{x^{2}-y^{2}}$ orbital, red squares indicate the $d_{xy}$ weight and green triangles indicate the $d_{3z^{2}-r^{2}}$ weight. The choice of valence band hopping parameters is listed in Appendix A, and we set $t_{c}=1,\Delta=15.5$. Symmetry guarantees that the overlaps of all $3d$ orbitals with the $s$-type conduction band orbital are the same, so we set $U_{\alpha}=U$ for all three valence orbitals. A short discussion of the effect of unequal $U_{\alpha}$ can be found in Appendix C.
Figure 4: Band structure for the 2D model. The conduction band is shown by the solid blue line. For the three valence bands, the blue circles indicate the weight of $d_{x^{2}-y^{2}}$ character, red squares indicate the $d_{xy}$ weight and green triangles indicate the $d_{3z^{2}-r^{2}}$ weight. The valence band hopping parameters are listed in Appendix A and $t_{c}=1,\Delta=15.5$. The band-gap is indirect, with the minimum along the $K-\Gamma$ direction, see next figure.
@@
Figure 5 shows the spectral weight $A_{\alpha\alpha}({\bf K},\omega)=-{1\over\pi}\text{Im}G_{\alpha\alpha}({\bf K},\bm{0},z)$ for $\bf K$} along high-symmetry lines in the Brillouin zone, for $\alpha=d_{x^{2}-y^{2}}$ (top panel), $\alpha=d_{xy}$ (middle panel) and $\alpha=d_{3z^{2}-r^{2}}$ (bottom panel). The thick black line marks the lower edge of the electron-hole continuum, $E_{\rm gap}({\bf K})$, with an indirect minimum gap lying nearly half-way between $K$ and $\Gamma$. Above this edge, the spectral function contains continuum states, but their weight is too small to be clearly visible on the color scale used. The relevant features are the three exciton bands visible well inside the gap (because of the large $U$ value used). These are discrete states, but here are broadened by the finite $\eta$ value we used. In the $d_{3z^{2}-r^{2}}$ projection (bottom panel) most of the intensity resides in the highest-energy (i.e. lowest binding energy) exciton branch. Along the $\Gamma-M-K$ cuts, the lowest energy exciton branch appears with comparable intensity in the top two panels, showing that it has comparable amounts of $d_{x^{2}-y^{2}}$ and $d_{xy}$ character for these momenta. By contrast, along the $K-\Gamma$ direction, the top panel shows two exciton branches while the middle panel shows a third, different branch. Clearly, along this cut there is no mixing between these two orbitals, and the excitons have either $d_{x^{2}-y^{2}}$ or $d_{xy}$ character.
@@
The simplest continuum approximation would predict that the lowest energy exciton appears along the $K-\Gamma$ line, at the momentum $\bf K$} for which $E_{\rm gap}({\bf K})$ has the global minimum. The middle panel indicates that the lowest energy exciton branch has indeed a local minimum at that momentum. However, a more careful inspection shows that the global minimum is actually at the $M$ point. We note that the $M$ point is not even a local minimum for $E_{\rm gap}({\bf K})$!
Figure 6: (top) Dispersion of the lowest-energy exciton $E_{\rm exc}({\bf K})$ for different values of $U$. The blue dashed line is the exciton energy calculated with a cut-off at $\delta_{M}=2$ (see Appendix G for details), and the dot marks its predicted lowest-energy exciton; (bottom) $E_{\rm exc}({\bf K})$ when ${\bf K}=M$ (blue triangles) and when $\bf K$} is the indirect gap, i.e. local minimum for $E_{\rm gap}({\bf K})$ along the $K-\Gamma$ line (red circles).
To understand what happens, we plot in the top panel of Fig. 6 the energy of the lowest-energy exciton branch along the same cuts, for increasing values of $U$. As expected, the exciton moves to lower energies as the attraction $U$ increases. Its dispersion closely follows the gap energy $E_{\rm gap}({\bf K})$ for small values of $U$, however, this changes as $U$ increases.
@@
Figure 8: Probability $|\phi_{\bf K}(\beta,\bm{\delta})|^{2}$ plotted as a function of the displacement $\bm{\delta}$ between electron and hole, for the lowest-energy exciton and for $\bf K$} equal to the indirect gap momentum, and $U=12$. The projections are on orbital $\beta=d_{x^{2}-y^{2}}$ (top), $\beta=d_{xy}$ (middle) and $\beta=d_{3z^{2}-r^{2}}$ (bottom). The triangular lattice sites are shown by the white circles. The probability is defined only at these lattice points; the interpolated continuous heat map is shown to guide the eye.
The bottom panel tracks the exciton energy at the momentum $M$ and at the indirect gap momentum lying on the $K-\Gamma$ line. The two curves cross at $U\approx 8$, marking a sharp transition in the momentum and character of the lowest-energy exciton. This is another example of a sharp transition that would be missed by a continuum approximation.
This change in the shape of the dispersion of the lowest-energy exciton band is illustrated graphically in Fig. 7, where its energy $E_{\rm exc}({\bf K})$ is plotted in the full Brillouin zone. For the smaller $U=2.5$ value (top panel) its minimum is clearly along the $\Gamma-K$ cut. However, as $U$ increases, the exciton energy at the $M$ point drops faster so that for $U=12$ (bottom panel) the minimum is at the $M$ point.
@@
This transition can be attributed to the symmetry already mentioned, which prevents mixing of $d_{x^{2}-y^{2}}$ and $d_{xy}$ orbitals along the $\Gamma-K$ line. This is supported by the plots of the real-space probability $|\phi_{\bf K}(\beta,\bm{\delta})|^{2}$ to find the hole in orbital $\beta$ and at a displacement $\bm{\delta}$ from the electron, shown in Fig. 8 when $\bf K$} is the indirect band-gap momentum, and in Fig. 9 when ${\bf K}=M$ point.
First, both of these excitons are Frenkel-like, with most of the probability concentrated at $\bm{\delta}=0$. The exceptions are the $d_{x^{2}-y^{2}}$ and the $d_{3z^{2}-r^{2}}$ orbitals (top and bottom panel in Fig. 8, respectively) where symmetry prevents weight at $\bm{\delta}=0$ and instead there is a very small amount of weight at $|{\bm{\delta}}|=a$. This is consistent with the vanishing spectral weight of this exciton in the top plot of Fig. 5. By contrast, the $M$ point exciton has considerable on-site probability in all orbitals, see Fig. (9). The reason why the M exciton lowers its energy faster with increasing $U$ is now clear: this exciton has the larger total probability of displacement $\bm{\delta}=0$ between the electron and hole, and therefore experiences more of the on-site $U$ attraction.
Such a discontinuous transition in the momentum of the lowest-energy exciton would again be completely missed by a continuum approximation. We note that the argument above can be made more quantitative by using a very simple variational approximation, where we allow this exciton to spread only over displacements $|\bm{\delta}|\leq 1$, i.e. we set $\delta_{M}=2$. The corresponding predicted exciton energy when $U=12$ is shown as the dashed blue line in Fig. 6(top). It is qualitatively correct, with the minimum energy (shown by the dot) in excellent agreement with the converged calculation, which has $\delta_{M}=10$. The quantitative differences are because at these $\bm{K}$} values, there is still some weight on states with $|{\bm{\delta}}|\geq 2$. Still, this shows that a very simple calculation involving only 7 sites (in this case) already captures the essential behavior of this small, Frenkel-like excitons. Details of this minimal calculation and further insights that can be obtained from it are presented in Appendix G.
@@ V Conclusions
In this work, we adopted a real-space approach to study exciton spectra and their wavefunctions in multi-orbital lattice models. The method can be used for excitons of any size, from Wannier-like to Frenkel-like. However, it becomes computationally most efficient as one approaches the limit of small, Frenkel-like excitons.
For the large Wannier excitons that appear in the limit of weak electron-hole attraction, we verified that our lattice results agree with the prediction of the corresponding continuum approximation. This comparison allowed us to propose a simple criterion to estimate when the continuum approximation is likely to fail quantitatively in models with a single valence and a single conduction band.
It is well known that continuum approximations must fail quantitatively, eventually, when applied to models that have small, Frenkel-like excitons because the full details of the bandstructure are relevant for these excitons' spectra. Here, we showed that the simplest continuum approximations can also fail qualitatively in multi-orbital models. Specifically, we presented two models where, in certain regions of the parameter space, the momentum of the lowest-energy exciton is not set by the momentum of the gap (whether the gap is direct or indirect). The simplest continuum approximations expand the valence/conduction bands about their maximum/minimum values, so by construction they anticipate that the lowest-energy exciton's momentum equals the momentum of the gap. More involved continuum approximations can be envisioned and these might be able to predict the qualitatively correct momentum, but we believe that for small, Frenkel-like excitons, it is easier to find the exact lattice spectrum with the method proposed here.
These examples are just beginning to explore possible new behavior of small, Frenkel-like excitons in models with multi-orbital bands. For such models, our real-space method provides a very efficient way to calculate the exciton spectra and their wavefunctions in the full Brillouin zone. Furthermore, the method can be extended to a few more carriers [33], relevant for calculating trions and spectra of larger complexes, to inclusion of couplings to bosonic degrees of freedom such as phonons or magnons, and to addition of interfaces and/or defects in the model. We believe that this makes it a valuable tool to explore various model Hamiltonians to understand possible behavior of small, Frenkel-like excitons.